Study/Discussion Questions on Collapse

(Note: These are adapted from some discussion questions on the publisher’s web site

for the book – http://us.penguingroup.com/static/rguides/us/collapse.html, and

a question concocted by Prof. Gareth Roberts.)

1. Exactly what does Diamond mean by a “collapse” and what are the five factors that Diamond identifies as contributing to the collapses he studies? (He uses these factors as a framework for analyzing each of the examples, so it will be helpful to have them in mind at all times.)

2. There is a lot of talk these days (especially from the more conservative end of the political spectrum) about how environmentalists have damaged their credibility by “crying wolf” – for example, issuing warnings about unchecked growth in human population and the effects of global warming that have not (yet) come true. Do you think Diamond is vulnerable to the charge of “crying wolf” in Collapse? If so, how would you dispute his analysis of the examples he cites? If not, why not?

3. “I am writing this book from a middle-of-the-road perspective,” writes Diamond in the introduction, “with experience of both environmental problems and of business realities” (p. 17). The middle of the road is often a tough place to be, since it opens one to attacks from both sides. How successful is Diamond in staking out this position? How does he balance (or fail to balance) environmental concerns with the needs of human societies?

4. “The past is a foreign country; they do things differently there,” is the famous opening line of a novel called The Go-Between by L. P. Hartley. How does Diamond’s point of view in Collapse relate to that attitude about the past? (For instance, would he say the past is not really that foreign after all, or would he argue that there is some key difference that sets us apart from previous civilizations and lessens the causes for concern about our ultimate fate?) What do you think about this question, after having read Collapse?

5. In Chapter 9, Diamond describes Tikopia as a kind of island paradise where natives saved their environment through eco-friendly gardening and devised a kind of rudimentary democratic system of government. Yet Tikopians also practiced infanticide and abortion to limit population growth. Were those practices immoral? Were they justified by the results? What does this say about our ability to judge the morality of past societies?

6. What view of human nature do you think underlies Collapse? Where do you think Diamond would stand on the nature vs. nurture debate, i.e., the role of genetics versus culture in determining human behaviors and responses?

7. How important are leaders in determining the success or failure (ecological or otherwise) of a civilization? To what extent did bad leadership contribute to or cause the collapses Diamond talks about? What about in our own culture? Do you think progress will come from enlightened leadership or rather from grassroots activism?

8. If the United States does collapse, how do you think it will happen? Which of Diamond’s five factors would play a role in the demise of American civilization as we know it? Do you think our collapse will occur suddenly, like the crash of Easter Island or the Classic Maya civilization, or is it more likely that we will experience a gradual but stable decline, as Great Britain did after World War II?

9. Which example of civilizational collapse described in the book do you find most compelling and why? Which best fits Diamond’s thesis? Diamond notes that “no other site that I have visited made such a ghostly impression on me as Rano Raraku, the quarry on Easter Island” (p. 79). Which image or passage in the book made the most powerful impression on you?

10. Diamond writes that our world “cannot sustain China and other Third World countries and current First World countries all operating at First World levels” (p. 376). (Comment – even though the first edition of Collapse came out in 2005, this is already quite out of date, of course, since China is well on its way to joining (leading?) the First World in 2011!) Can we ethically deny Third World countries the comforts and advantages that we in the First World enjoy? In your opinion, what should we do to lessen or resolve looming conflicts over resources between First and Third World countries? Or is there “enough to go around” if it is allocated rationally (e.g. by market forces)?

11. Diamond reveals that while writing the book he found himself lurching between hope and despair. What emotions did Collapse inspire in you? Did you come away depressed, cautiously hopeful, or did you have an entirely different reaction?

12. Part of the mission of Holy Cross is to foster a community which asks and discusses fundamental questions such as “What are our obligations to one another?” and “What is our special responsibility to the world’s poor and powerless?” With that in mind, what personal obligation or responsibility, if any, do you have to address the challenges posed in Diamonds work, particularly those raised in the final chapters? Are there particular strengths, experiences, or talents you possess that would be helpful in confronting some of these issues? Has this book encouraged you to consider any changes in your future educational or vocational path? Are there any courses, service opportunities, or travel that you would like to pursue during your Holy Cross education related to these issues?